A bicycle has a momentum of 24 kg m/s. What momentum would the bicycle have if it had…


… twice the mass and was moving at the same speed?

Answer 1


… the same mass and was moving with one-half the speed?

Answer 2


… three times the mass and was moving with twice the speed?

Answer 3


… one-half the mass and was moving with twice the speed?

Answer 4


… the same mass and was moving with twice the speed?

Answer 5


… three times the mass and was moving with one-half the speed?


1
Expert's answer
2022-04-15T16:40:16-0400

The momentum equals:

"p=mv."


If the bicycle had twice the mass and was moving at the same speed:

"m'=2m,\space v'=v."

"p'=m'v'=2mv=2p=2\cdot 24=48\space kg\cdot m/s."

Answer 1: "48\space kg\cdot m/s."


If the bicycle had the same mass and was moving with one-half the speed?

"m'=m,\space v'=v/2."

"p'=m\cdot v/2=p/2=24/2=12\space kg\cdot m/s."

Answer 2: "12\space kg\cdot m/s."


If the bicycle had three times the mass and was moving with twice the speed?

"m'=3m,\space v'=2v."

"p'=3m\cdot 2v=6mv=6p=6\cdot 24=144\space kg\cdot m/s."

Answer 3: "144\space kg\cdot m/s."


If the bicycle had one-half the mass and was moving with twice the speed?

"m'=m/2,\space v'=2v."

"p'=(m/2)\cdot 2v=mv=p=24 \space kg \cdot m/s."

Answer 4: "24 \space kg \cdot m/s."


If the bicycle had the same mass and was moving with twice the speed?

"m'=m,\space v'=2v,"

"p'=m\cdot 2v=2mv=2p=2\cdot 24=48 \space kg\cdot m/s."

Answer 5: "48 \space kg\cdot m/s"


If the bicycle had three times the mass and was moving with one-half the speed?

"m'=3m,\space v'=v/2."

"p'=3m\cdot v/2=(3/2)mv=(3/2)p=(3/2)\cdot 24=36\space kg\cdot m/s".

Answer 6: "36\space kg\cdot m/s"